OpenStudy (anonymous):

How would I solve : find the point on the line 2x-4y = 3 that is closest to hte origin using lagrange multipliers????

OpenStudy (anonymous):

let a point on the line (a,b) satisfy it with equation of line- 2a-4b=3---------equation 1 now slope of the line joining (a,b) to origin is b/a,and slope of given line is 1/2 now (1/2)X(b/a)=-1 by simplifying 2a+b=0------------equation 2 by solving b=-3/5 and a=3/10

OpenStudy (anonymous):

Wondelful! Thanks

OpenStudy (anonymous):

welcome.....